A beverage wholesaler wants to create a new punch. He will mix fruite juice worth $3 a gallon and rum worth $8 a gallon. He wants to obtain 145 gallons worth of punch worth $5 a gallon. How much of each beverage should he use?
Answer by ptaylor(2198)

OK
Let x=amount of rum needed
Then 145-x=amount of fruit juice needed
Now we know that the total value of the rum and fruit juice before they are mixed together is equal to the total value of the final mixture.
So: ($ understood)
8x+3(145-x)=5*145
8x+435-3x=725 subtract 435 from each side and collect like terms
5x=290
x=58 gal-----------------------amount of rum needed
145-58=87 gal----amount of fruit juice needed
CK
8*58+3*87=5*145
464+261=725
725=725
Hope this helps--ptaylor